RCT working

The mirror support compressor’s pressure switch, which activates and deactivates the pump, burned out. It was replaced with a new switch, one with a bit more control on its activation pressure, which should reduce the wear on this part.

The problem with the mirror support compressor seems to be resolved, and the telescope should be working normally. I will check the focus and pointing again tonight, then resume the scheduler.

Leave a Reply

Your email address will not be published.

You may use these HTML tags and attributes: <a href="" title=""> <abbr title=""> <acronym title=""> <b> <blockquote cite=""> <cite> <code> <del datetime=""> <em> <i> <q cite=""> <strike> <strong>